OpenStudy (anonymous):

Solve 3x^2-5x-7=0 What are the solutions? x=

OpenStudy (akshay_budhkar):

Do you know the quadratic formula?

OpenStudy (anonymous):

A little

OpenStudy (akshay_budhkar):

do you know x=\[(-b+-\sqrt{b^2-4ac})/2a\]

OpenStudy (akshay_budhkar):

this is the quadratic formula

OpenStudy (anonymous):

yeah, but to me it is confusing

OpenStudy (akshay_budhkar):

it is a formula for ax^2+bx+c=0.. ok?

OpenStudy (anonymous):

ok

OpenStudy (akshay_budhkar):

now you compare the terms of your equation with it. to get a,b and c

OpenStudy (akshay_budhkar):

a=3,b=-5,c=-7

OpenStudy (akshay_budhkar):

now just subsitute in the formula..

OpenStudy (anonymous):

how do you find out what a, b, and c are?

OpenStudy (akshay_budhkar):

ok

OpenStudy (anonymous):

and what do I do with x?

OpenStudy (akshay_budhkar):

see ax^2+bx+c=0 ......3x^2-5x-7=0.. compare the two

OpenStudy (akshay_budhkar):

a=3 b=-5 c=-7

OpenStudy (anonymous):

oops I see what you mean about a b and c.

OpenStudy (akshay_budhkar):

lol.. so you get it now? Now simply subsitute it in the formula.:)

OpenStudy (anonymous):

I believe so. I hope so

OpenStudy (akshay_budhkar):

you can view this for final answer and the graph.. http://www.wolframalpha.com/input/?i=3x^2-5x-7%3D0++

OpenStudy (akshay_budhkar):

Hope it helps.. if you any more doubt feel free to ask:) n dun forget my medal..lol

OpenStudy (anonymous):

thank you